A container at an initial volume of 0.07 m^3 holds
Answer #1

In adiabatic process

PV^(alpha) = constant

alpha = 1.66 for an ideal monoatomic gas

so,

P1V1^(alpha) = P2V2^(alpha)

1*(0.07/0/03)^1.66 = P2 = 4.08 atm

V2T2^(1/1.66-1) = V1 T1(1/1.66-1)

T2 = 611.9 k